# Spatial-Temporal Decoupling Contrastive Learning for Skeleton-based Human Action Recognition

### v3

@@ Abstract
Skeleton-based action recognition is a central task in human-computer interaction. However, most previous methods suffer from two issues: (i) semantic ambiguity arising from spatial-temporal information mixture; and (ii) overlooking the explicit exploitation of the latent data distributions (i.e., the intra-class variations and inter-class relations), thereby leading to sub-optimum solutions of the skeleton encoders. To mitigate this, we propose a spatial-temporal decoupling contrastive learning (STD-CL) framework to obtain discriminative and semantically distinct representations from the sequences, which can be incorporated into various previous skeleton encoders and can be removed when testing. Specifically, we decouple the global features into spatial-specific and temporal-specific features to reduce the spatial-temporal coupling of features. Furthermore, to explicitly exploit the latent data distributions, we employ the attentive features to contrastive learning, which models the cross-sequence semantic relations by pulling together the features from the positive pairs and pushing away the negative pairs. Extensive experiments show that STD-CL with four various skeleton encoders (HCN, 2S-AGCN, CTR-GCN, and Hyperformer) achieves solid improvements on NTU60, NTU120, and NW-UCLA benchmarks. The code will be released at https://github.com/LibertyZsj/STD-CL soon.
@@ 1 Introduction
Human action recognition is an active task in the computer vision area. Due to its wide range of applications in human-computer interaction, video analysis, virtual reality, and so on, this task has been researched extensively in the past decade Ren et al. (2020). In recent years, with the development of depth sensors Cao et al. (2017) and human pose estimation algorithms Dang et al. (2019), the skeleton sequence consisting of coordinates of human joints can be easily acquired. The skeleton-based action recognition utilizing human joints has attracted much interest due to its robustness to background clutter, illumination, and viewpoint changes Zhang et al. (2019).
Figure 1: The framework of action recognition in the previous methods. A skeleton encoder extracts the spatial-temporal co-occurrence feature, and a softmax-based linear classifier is used to project the feature to the category distributions.
Figure 2: Visualization of samples from two confused classes (“put on a hat” (blue) and “put on glasses” (red). (a) Visualization of sequences performed by different persons (each row represents a different person) in the same view. (b) Visualization of feature representation by t-SNE for the corresponding action in the test set. The left is from the CTR-GCN, while the right is from our method. Each color denotes a certain class.
@@
Generally, most of the previous approaches Shi et al. (2019); Chen et al. (2021a, b) in skeleton-based action recognition follow a paradigm as shown in Figure 1. A skeleton encoder aggregates the spatial-temporal co-occurrence feature Li et al. (2018), and a softmax-based linear classifier is used to project the feature to the category distributions. However, this paradigm has some limitations. The main problems are semantic ambiguity arising from the spatial-temporal information mixture in the co-occurrence feature and the absence of explicit exploitation of the latent data distributions.
The movements of joints in the skeleton sequences can be decoupled into spatial dimension and temporal dimension. The spatial dimension indicates the topological relations and the temporal dimension indicates the trajectory of action, each reflecting distinct semantics at the corresponding dimension Xu et al. (2023). For example, temporal-specific information is more important when distinguishing the actions “put on a shoe” and “take off a shoe”, which differ more in the frame-level dimension. Conversely, the actions “point to something” and “throw”, have more spatial differences in the joint-level dimension, so spatial-specific information is more important to distinguish them. The skeleton encoders blend the spatial-temporal information in their aggregations, leading to a mixture of spatial-specific and temporal-specific features and indistinctiveness of semantics in the global features. Thus, the discriminative capability of existing skeleton encoders is limited by the semantic ambiguity arising from the spatial-temporal information mixture.
@@
Nevertheless, as shown in Figure 2, both the action “put on a hat” and the action “put on glasses” contains a crucial sub-action “lift arms”, leading to semantic similarity and less distinct inter-class distances to some extent. Furthermore, due to individual differences and the unknowability of the external environment, actions within the same category exhibit intra-class variation. The linear classifier is typically trained to optimize accuracy only, paying less attention to exploiting the latent data distributions Wang et al. (2023). The t-SNE Van der Maaten and Hinton (2008) distribution of features extracted from CTR-GCN Chen et al. (2021a) is shown in Figure 2 (b), which can’t locate a distinguishable classification boundary. Therefore, the absence of explicit modeling of the intra-class variation and inter-class relation is insufficient for recognizing the ambiguous samples.
@@
To alleviate these drawbacks, we propose a spatial-temporal decoupling contrastive learning (STD-CL) framework to encourage the skeleton encoders to learn more discriminative and semantically distinct representations. It’s worth noting that our STD-CL could be combined with previous skeleton encoders and can be removed naturally without additional consumption when testing. In the training phase, the global features are decoupled into spatial attentive features and temporal attentive features through the designed spatial-temporal feature decoupling (STFD) module. Moreover, we build two memory banks to store the decoupled spatial and temporal features respectively, and sample the positives and negatives from them according to labels. In particular, we employ contrastive learning to model the cross-sequence semantic relations by pulling together the decoupled features from positive pairs and pushing away the features from negative pairs. With constant contrast, the skeleton encoders are forced to explicitly explore the latent data distributions with distinct inter-class relations and reduced intra-class variations. Meanwhile, since the features used for contrast are decoupled, the degree of spatial-temporal information coupling within global features decreases so that the semantics become more explicit. Consequently, in the testing phase, the skeleton encoders can directly predict the categories with higher accuracy.
@@
  • To extract more discriminative and semantically distinct features from the skeleton sequences, we proposed a novel STD-CL framework to decouple the features into spatial- and temporal-specific features and apply them for contrastive learning to explore the latent data distributions.
  • The proposed STD-CL can be seamlessly incorporated into various previous skeleton encoders, which can be regarded as plug-and-play in the training stage and can be removed at the testing stage.
  • Extensive experiments show that STD-CL achieves significant improvements based on four various state-of-the-art methods (HCN Li et al. (2018), 2S-AGCN Shi et al. (2019), CTR-GCN Chen et al. (2021a), and Hyperformer Zhou et al. (2022)) on three benchmark datasets (NTU60, NTU120, and NW-UCLA).
Figure 3: Overview of the proposed methods. The black dotted lines represent the training process and the red solid line represents the testing process. (a) The framework of our proposed STD-CL. The skeleton encoder extracts spatial-temporal feature $\boldsymbol{X}$ from the input skeleton sequence $\boldsymbol{S}$. Then the feature $\boldsymbol{X}$ is fed into the STFD module to decouple spatial attentive feature $\boldsymbol{s}$ and temporal attentive feature $\boldsymbol{t}$. Two memory banks are built to store the decoupled features respectively. We contrast the positive and negative instance features sampled from the memory banks according to the labels. The memory banks and the contrastive learning process are only used in the training phase and have no impact in the testing phase. (b) The architecture of the STFD module.
@@ 2.1 Skeleton-based Action Recognition
Skeleton-based action recognition aims to classify actions from sequences of human keypoints. In early studies, RNN was a natural choice to handle the skeleton sequences. HBRNN Du et al. (2015) applied a hierarchical RNN architecture to model the long-range temporal dependency of the skeleton sequences. STA-LSTM Song et al. (2017) proposed an end-to-end spatial and temporal attention model to focus on discriminative joints of the skeleton within each frame of the inputs. Inspired by the success of CNN in image recognition, CNN-based methods have also been studied. HCN Li et al. (2018) treated the joint dimension as channels to aggregate different levels of spatial-temporal context information. TA-CNN Xu et al. (2022) adopted a pure CNN architecture to model the irregular skeleton topology.
@@
The human body can be abstracted as a graph structure, which treats the joint as a node and the bone between two joints as an edge. Thus, GCN-based methods were introduced to this task. ST-GCN Yan et al. (2018) adopted GCN on the predefined spatial-temporal graphs to model the relations between joints. 2S-AGCN Shi et al. (2019) modeled the correlation between two joints given corresponding features with the self-attention mechanism. CTR-GCN Chen et al. (2021a) proposed a channel-wise topology refinement graph convolution to model fine-grained relation. InfoGCN Chi et al. (2022) combined a learning objective and an encoding method to break the information bottleneck.
@@
With the popularity of transformers in computer vision, transformer-based methods have also been investigated for this task. ST-TR Plizzari et al. (2021) proposed a two-stream transformer architecture to model the spatial and temporal dimensions respectively. Hyperformer Zhou et al. (2022) applied transformer to model spatial-temporal features from the hyper-graph of the sequences.
2.2 Contrastive Learning
@@
Recently, contrastive learning has been widely used in self-supervised representation learning. The core idea is to learn the discriminative features by pulling the positive pairs and pushing away the negative pairs in the embedded space. The positive pairs are features from transformed sample versions with different augmentations, and the negative pairs are from different samples. InstDisc Wu et al. (2018) maintained a memory bank for storing representations. SimCLR Chen et al. (2020) proposed to adopt a huge batch size to learn more useful representations. MoCo He et al. (2020) built a dynamic dictionary with a momentum update based on a memory bank to keep the stored representations consistent. Inspired by this, our work tries to introduce memory bank mechanisms for contrastive learning to obtain discriminative features.
@@
Contrastive learning was usually adopted in prior works Xu et al. (2023) to learn the invariance from the skeleton sequences for unsupervised learning in this field. Moreover, several works are applying contrastive learning to supervised action recognition. GAP Xiang et al. (2023) proposed a framework for contrasting the features from visual and text modalities to enhance the representation by using knowledge about actions and human body parts. SkeletonGCL Huang et al. (2023) explicitly explored the rich cross-sequence relations by using graph contrastive learning. However, SkeletonGCL can only be incorporated into the GCN-based methods, which limits its generalization. FR-Head Zhou et al. (2023) employed multi-level contrastive learning in GCN-based methods to distinguish ambiguous actions.
However, the methods mentioned above are more like surgery for GCN-based skeleton encoders, leading to limitations on generalizability to other types of methods.
@@ 3 Methodology
3.1 Perliminary
@@
Skeleton-based action recognition is to predict the categories from the input sequences. We define the human skeleton sequence with $J$ keypoints and $T_{0}$ frames in 3D space as $\boldsymbol{S}\in\mathbb{R}^{J\times T_{0}\times 3}$. The sequence is fed into a skeleton encoder $\mathcal{F}$ to extract the features $\boldsymbol{X}\in\mathbb{R}^{J\times T\times C}$. The skeleton encoder aggregates spatial and temporal information to obtain discriminative spatial-temporal features. After fully spatial-temporal aggregations, a global pooling (GP) layer is adapted to summarize the global features. Finally, a softmax-based fully-connected (FC) layer maps the global feature to a probability prediction $\hat{\boldsymbol{y}}\in\mathbb{R}^{K}$ of $K$ candidate categories. The process can be defined as:
$\boldsymbol{X}=\mathcal{F}(\boldsymbol{S})$ (1)
$\boldsymbol{y}=FC(GP(\boldsymbol{X}))$ (2)
@@
In the process of optimization, a standard cross-entropy loss is applied to supervise the prediction distribution and the ground truth $\boldsymbol{y}$ as follows:
$\mathcal{L}_{CE}=-\sum_{i}\boldsymbol{y}_{i}log\hat{\boldsymbol{y}_{i}}$ (3)
[…]
Dataset NTU 60 NTU 120
Setting X-Sub X-View X-Sub X-Set
Method/Streams J B 2S 4S J B 2S 4S J B 2S 4S J B 2S 4S
VA-LSTM Zhang et al. (2017) 79.4* 87.6* -
AGC-LSTM Si et al. (2019) - - 89.2* - - - 95.0* - - - - - - - - -
ST-GCN Yan et al. (2018) 81.5 - - - 88.3 - - - - - - - - - - -
SGCN Zhang et al. (2020) - - 89.0 - - - 94.5 - - - 79.2 - - - 81.5 -
Shift-GCN Cheng et al. (2020b) 87.8 - 89.7 90.7 95.1 - 96.0 96.5 80.9 - 85.3 85.9 83.2 - 86.6 87.6
DC-GCN+ADG Cheng et al. (2020a) - - 90.8 - - - 96.6 - - - 86.5 - - - 88.1 -
Dynamic GCN Ye et al. (2020) - - - 91.5 - - - 96.0 - - - 87.3 - - - 88.6
MS-G3D Liu et al. (2020) 89.4 90.1 91.5 - 95.0 95.3 96.2 - - - 86.9 - - - 88.4 -
DSTA Shi et al. (2020) - - 91.5 - - - 96.4 - - - 86.6 - - - 88.0 -
MST-GCN Chen et al. (2021b) 89.0 89.5 91.1 91.5 95.1 95.2 96.4 96.6 82.8 84.8 87.0 87.5 84.5 86.3 88.3 88.8
ST-TR Plizzari et al. (2021) - - 89.9 - - - 96.1 - - - 82.7 - - - 84.7 -
TA-CNN Xu et al. (2022) 88.9 89.2 91.0 91.5 94.5 94.1 95.7 95.9 84.0 85.1 87.8 88.2 85.3 86.3 89.0 89.6
EfficientGCN-B4 Song et al. (2022) - - 91.7 - - - 95.7 - - - 88.3 - - - 89.1 -
InfoGCN Chi et al. (2022) 89.4 90.6 91.3 92.3 95.2 95.4 96.2 96.7 84.2 86.9 88.2 89.2 86.3 88.5 89.4 90.7
HCN Li et al. (2018) - - 84.3* - - - 89.9* - - - - - - - - -
HCN w/ STD-CL - - 85.1* - - - 91.2* - - - - - - - - -
2S-AGCN Shi et al. (2019) 88.9 89.2 91.0 91.5 94.5 94.1 95.7 95.9 83.8 84.9 87.7 88.1 85.3 86.3 89.0 89.6
2S-AGCN w/STD-CL 89.4 89.8 91.5 91.9 94.7 94.6 96.1 96.3 84.1 85.1 87.9 88.2 86.0 86.6 89.4 89.9
CTR-GCN Chen et al. (2021a) 89.8 90.2 92.0 92.4 94.8 94.8 96.3 96.8 84.9 85.7 88.7 88.9 86.7 87.5 90.1 90.5
CTR-GCN w/STD-CL 90.4 90.8 92.2 92.7 95.2 95.0 96.4 96.9 85.6 86.7 89.3 89.5 86.8 88.2 90.4 90.9
Hyperformer Zhou et al. (2022) 90.3 91.1 92.0 92.7 94.5 94.4 95.5 96.2 86.1 87.4 88.9 89.9 87.8 89.0 90.6 91.2
Hyperformer w/STD-CL 90.8 91.3 92.3 92.9 95.0 94.7 95.9 96.4 86.8 88.1 89.2 90.1 88.3 89.2 90.7 91.4
3.2 Spatial-Temporal Decoupling Contrastive Learning
Human actions have the property of spatio-temporal coupling, the feature $\boldsymbol{X}$ after aggregation contains complex spatial and temporal relations, which may mix the spatial and temporal information. Thus, we decouple the feature into the spatial-aware and temporal-aware features to reduce the spatial-temporal information coupling. To further obtain more discriminative features, we apply the decoupled features to contrastive learning to explore the cross-sequence semantic relations. The core is to explore the latent data distributions by pulling together the features from the positive pairs and pushing away the negative pairs.
3.2.1 Spatial Temporal Feature Decoupling

### v4

# A Generically Contrastive Spatiotemporal Representation Enhancement for 3D Skeleton Action Recognition

@@ Abstract
Skeleton-based action recognition is a central task in computer vision and human-robot interaction. However, most previous methods suffer from overlooking the explicit exploitation of the latent data distributions (i.e., the intra-class variations and inter-class relations), thereby leading to confusion about ambiguous samples and sub-optimum solutions of the skeleton encoders. To mitigate this, we propose a Contrastive Spatiotemporal Representation Enhancement (CSRE) framework to obtain more discriminative representations from the sequences, which can be incorporated into various previous skeleton encoders and can be removed when testing. Specifically, we decompose the representation into spatial- and temporal-specific features to explore fine-grained motion patterns along the corresponding dimensions. Furthermore, to explicitly exploit the latent data distributions, we employ the attentive features to contrastive learning, which models the cross-sequence semantic relations by pulling together the features from the positive pairs and pushing away the negative pairs. Extensive experiments show that CSRE with five various skeleton encoders (HCN, 2S-AGCN, CTR-GCN, Hyperformer, and BlockGCN) achieves solid improvements on five benchmarks. The code will be released at https://github.com/zhshj0110/CSRE.
keywords:
human-robot interaction, action recognition, skeleton, contrastive learning
journal: Journal of Pattern Recognition
@@ 1 Introduction
Human action recognition is an active task in the computer vision area. Due to its wide range of applications in human-computer interaction, video analysis, virtual reality, and so on, this task has been researched extensively in the past decade [1]. In recent years, with the development of depth sensors [2] and human pose estimation algorithms [3], the skeleton sequence consisting of coordinates of human joints can be easily acquired. Skeleton-based action recognition utilizing human joints has attracted much interest due to its robustness to background clutter, illumination, and viewpoint changes [1].
@@
Generally, most of the previous approaches [4, 5, 6] in skeleton-based action recognition follow a framework as shown in Figure 1. A skeleton encoder aggregates the spatiotemporal co-occurrence feature [7], and a softmax-based linear classifier is used to project the pooling feature to the category distributions. However, this widespread framework has some limitations. The main problem is the absence of explicit exploitation of the latent data distributions, thereby leading to confusion about ambiguous samples and sub-optimum solutions of skeleton encoders.
@@
Specifically, because the important interactive objects and contexts are omitted in skeletonization, it is difficult to distinguish the action categories with similar motion patterns significantly. For example, as shown in Figure 2 (a), the action “take off glasses" and the action “take off a hat" exhibit a high degree of similarity due to the absence of the interactive objects (the hat and glasses). Furthermore, due to individual differences and the unknowability of the external environment, actions within the same category exhibit a wide range of intra-class variation. Although previous skeleton encoders extract discriminative representations from the skeleton sequences for classification, the linear classifier is typically trained to optimize accuracy only, paying less attention to exploiting the latent data distributions [8]. The t-SNE [9] visualization of features extracted from CTR-GCN [5] is shown in Figure 2 (b), which can’t locate a distinguishable classification boundary because of the ambiguous samples. Therefore, we suppose explicit modeling of the intra-class variation and inter-class relation is crucial for recognizing the ambiguous samples.
Figure 1: Comparison between previous method (in green) and proposed CSRE framework (in blue). Compared to the previous method, our CSRE can enable the skeleton encoder better to explore the fine-grained motion patterns along the spatial and temporal dimensions, which is beneficial for recognizing ambiguous action samples.
@@
To alleviate this drawback, As shown in 1, we propose a generically Contrastive Spatiotemporal Representation Enhancement (CSRE) framework to encourage the skeleton encoders to learn more discriminative representations to distinguish ambiguous samples. It’s worth noting that our CSRE can be combined with previous skeleton encoders and removed naturally without additional consumption when testing. In the training phase, to enable the network to better explore the fine-grained motion patterns among ambiguous actions along the topological and temporal dimensions, the spatiotemporal representation is decomposed into spatial and temporal components using the designed spatial-temporal feature decomposition (STFD) module. Moreover, we build two memory banks to store the spatial-specific and temporal-specific features, respectively, and sample the positives and negatives from them according to labels. In particular, we employ contrastive learning to model the cross-sequence semantic relations by pulling together the decomposed features from positive pairs and pushing away the features from negative pairs. With constant contrast, the skeleton encoders are forced to explicitly explore the latent data distributions with distinct inter-class relations and reduced intra-class variations. Consequently, in the testing phase, the skeleton encoders can directly predict the categories with higher accuracy.
Figure 2: Visualization of samples from two confused classes (“take off glasses" (blue) and “take off a hat" (red). (a) Visualization of sequences performed by different persons (each row represents a different person) in the same view. (b) Visualization of feature representation by t-SNE for the corresponding action in the test set. The left is from the CTR-GCN, while the right is from our method. Each color denotes a certain class.
@@
  • To extract more discriminative representation to distinguish the ambiguous samples, we proposed a novel and generically CSRE framework to decompose the features into spatial- and temporal-specific features and apply them for contrastive learning to explore the latent data distributions explicitly.
  • The proposed CSRE can be seamlessly incorporated into various previous skeleton encoders, which can be regarded as plug-and-play in the training stage and can be removed at the testing stage.
  • Extensive experiments show that CSRE achieves significant improvements based on five various state-of-the-art methods (HCN [7], 2S-AGCN [4], CTR-GCN [5], Hyperformer [10], and BlockGCN [11]) on five benchmarks.
@@ 2.1 Skeleton-based Action Recognition
Skeleton-based action recognition aims to classify actions from sequences of human keypoints. In early studies, Recurrent neural network (RNN) based methods [12, 13] and Convolution neural network (CNN) based methods [7, 14, 15] are popular choices for solving the skeleton-based action recognition problem. However, both CNN-based and RNN-based methods overlook human topology and the spatial interactions between joints.
@@
The human body can be abstracted as a graph structure, which treats the joint as a node and the bone between two joints as an edge. Thus, GCN-based methods were introduced to this task. ST-GCN [16] adopted GCN on the predefined spatial-temporal graphs to model the relations between joints. 2S-AGCN [4] modeled the correlation between two joints given corresponding features with the self-attention mechanism. CTR-GCN [5] proposed a channel-wise topology refinement graph convolution to model fine-grained relations. InfoGCN [17] combined a learning objective and an encoding method to break the information bottleneck. HD-GCN [18] effectively constructs an HD-Graph by decomposing every joint node into several sets to extract major adjacent and distant edges. DD-GCN [19] constructs the directed diffusion graph for action modeling and introduces the activity partition strategy. BlockGCN [11] introduces BlockGC to streamline multi-relational modeling, reducing parameters while improving performance over traditional GCNs.
@@
With the popularity of transformers in computer vision, transformer-based methods have also been investigated for this task. ST-TR [20] proposed a two-stream transformer architecture to model the spatial and temporal dimensions, respectively. Hyperformer [10] applied transformer to model spatial-temporal features from the hyper-graph of the sequences. ISTA-Net [21] introduces the transformer to model spatial, temporal, and interactive relations simultaneously. STC-Net [22] proposes the novel STC module and Dilated Kernels for Graph Convolution, enabling adaptive spatiotemporal coverage and long-range dependency modeling without increasing parameters. SelfGCN [23] combines graph convolution and self-attention via the MFSG module for modeling both local and global joint relationships and the TSSA module for learning temporal-specific spatial features, addressing limitations in long-range dependency modeling and uniform skeleton topology.
2.2 Spatiotemporal Feature Decomposition
Human action has the property of spatial-temporal coupling. Thus, previous methods [24, 20, 25] tried to decompose features into specific dimensions and utilized the dual-branch or sequential network architecture to extract temporal-spatial features. STA [24] proposed a spatiotemporal attention module that simultaneously learns frame-level and channel-level attention weights to enhance 3D CNNs’ ability to capture discriminative spatial and temporal features for video-based action recognition. ST-TR [20] proposed a novel two-stream Transformer-based model that employs Spatial Self-Attention (SSA) and Temporal Self-Attention (TSA) modules to dynamically model spatial and temporal dependencies in skeleton data, thereby effectively decomposing and capturing the spatial and temporal features for action recognition. RSA-Net [25] proposed a Relation-mining Self-Attention Network that decomposes spatiotemporal features by using pairwise self-attention to capture joint relationships and unary self-attention to model global dependencies while also separately computing action semantics and positional information with distinct projection matrices. Although these previous methods have attempted to decompose features into spatial and temporal-specific features, they still lack the explicit modeling of the latent data distributions along the specific dimensions.
2.3 Contrastive Learning in Skeleton-based Action Recognition
@@
Recently, contrastive learning has been widely used in self-supervised representation learning. The core idea is to learn the discriminative features by pulling the positive pairs and pushing away the negative pairs in the embedded space. The positive pairs are features from transformed sample versions with different augmentations, and the negative pairs are from different samples. InstDisc [26] maintained a memory bank for storing representations. SimCLR [27] proposed adopting a huge batch size to learn more useful representations. MoCo [28] built a dynamic dictionary with a momentum update based on a memory bank to keep the stored representations consistent. Inspired by this, our work tries to introduce memory bank mechanisms for contrastive learning to obtain discriminative features.
@@
Contrastive learning was usually adopted in prior works [29, 30, 31] to learn the invariance from the skeleton sequences for unsupervised learning in this field. Moreover, several works are applying contrastive learning to supervised action recognition. GAP [32] proposed a framework for contrasting the features from visual and text modalities to enhance the representation by using knowledge about actions and human body parts. SkeletonGCL [33] explicitly explored the rich cross-sequence relations by using graph contrastive learning. However, SkeletonGCL can only be incorporated into the GCN-based methods, which limits its generalization. FR-Head [34] employed multi-level contrastive learning in GCN-based methods to distinguish ambiguous actions. However, the methods mentioned above are more like surgery for GCN-based skeleton encoders, leading to limitations on generalizability to other types of methods.
@@ 3 Methodology
In this section, we elaborate on the structure and training methodology of CSRE. Firstly, we state the problem formulation of skeleton-based action recognition and the generic modeling progress of previous approaches in Section 3.1. Subsequently, we introduce the structure of the proposed STFD module in Section 3.2. Finally, we present the overall framework and training objective in Section 3.3.
3.1 Preliminaries
@@
Skeleton-based action recognition is to predict the categories from the input sequences. We define the human skeleton sequence with $J$ keypoints and $T_{0}$ frames in 3D space as $\bm{S}\in\mathbb{R}^{J\times T_{0}\times 3}$. The sequence is fed into a skeleton encoder $\mathcal{F}$ to extract the features $\bm{X}\in\mathbb{R}^{J\times T\times C}$. The skeleton encoder aggregates spatial and temporal information to obtain discriminative spatial-temporal features. After fully spatial-temporal aggregations, a global pooling (GP) layer is adapted to summarize the global features. Finally, a softmax-based fully-connected (FC) layer maps the global feature to a probability prediction $\hat{\bm{y}}\in\mathbb{R}^{K}$ of $K$ candidate categories. The process can be defined as:
$\bm{X}=\mathcal{F}(\bm{S})$ (1)
$\bm{y}=FC(GP(\bm{X}))$ (2)
@@
In the process of optimization, a standard cross-entropy loss is applied to supervise the prediction distribution and the ground truth $\bm{y}$ as follows:
$\mathcal{L}_{CE}=-\sum_{i}\bm{y}_{i}log\hat{\bm{y}_{i}}$ (3)
3.2 Spatial Temporal Feature Decomposition
Figure 3: The architecture of the STFD module.
The skeleton sequences of human actions have the property of spatiotemporal coupling, and each dimension in the encoded representation $\bm{X}$ reflects distinct semantic information: the spatial dimension contains the topological relations, the temporal dimension includes the motion trajectories, and the channel dimension encompasses the richness of features. To explore fine-grained motion patterns along corresponding dimensions, we decompose the representation into the spatial-aware and temporal-aware features.